solve by substitution
8x+4y=4
9x-4y=-38

Answer by mananth(16946)

8x+4y=4 .............1
rewrite the second equation
4y=38+9x---------2
/4
y=9.5+2.25x
substitute the value of y in (1)
8x+4(9.5+2.25x)=4
8x+38+9x=4
8x+9x=4-38
17x=-34
/17
x=-2
Plug the value of x in (1)
8x+4y=4
8*-2+4y=4
-16+4y=4
4y=20
/4
y=5